# PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection

### v3

@@ 3.5 Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both.
The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))$ (7)
$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$
$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))$ (8)
$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
@@
The final PatchAD loss consists of the aforementioned two components, with a constraint coefficient used to control the strength of the constraint:
$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}.$ (11)
@@ 3.6 Anomaly Score
${y}_{i}=\left\{\begin{aligned} &1\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})\geq\sigma\\ &0\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})<\sigma.\end{aligned}\right.$ (13)
Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best ones are in Bold, and the second ones are underlined.
@@
OCSVM Schölkopf et al. (2001) 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
IForest Liu et al. (2008) 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
LOF Breunig et al. (2000) 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
U-Time Perslev et al. (2019) 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
ITAD Shin et al. (2020) 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
DAGMM Zong et al. (2018) 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
VAR Anderson (1976) 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD Yairi et al. (2017) 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA Tariq et al. (2019) 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 Deldari et al. (2021) 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
BeatGAN Zhou et al. (2019) 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM-VAE Park et al. (2018) 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BOCPD Adams and MacKay (2007) 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
Deep-SVDD Ruff et al. (2018) 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
LSTM Hundman et al. (2018) 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly Su et al. (2019) 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion Li et al. (2021b) 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC Shen et al. (2020) 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans Xu et al. (2021) 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetector Yang et al. (2023) 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
@@
Table 2: Overall results on NIPS-TS datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
Metric P R F1 P R F1
MatrixProfile Yeh et al. (2016) 16.70 17.50 17.10 4.60 18.50 7.40 12.25
GBRT Elsayed et al. (2021) 44.70 37.50 40.80 17.50 14.00 15.60 28.20
LSTM-RNN Bontemps et al. (2016) 52.70 22.10 31.20 34.30 27.50 30.50 30.85
Autoregression Rousseeuw and Leroy (2005) 42.10 35.40 38.50 39.20 31.40 34.90 36.70
OCSVM Schölkopf et al. (2001) 47.40 49.80 48.50 18.50 74.30 29.60 39.05
IForest* Liu et al. (2008) 40.60 42.50 41.60 39.20 31.50 39.00 40.30
AutoEncoder Sakurada and Yairi (2014) 49.70 52.20 50.90 42.40 34.00 37.70 44.30
AnomalyTrans Xu et al. (2021) 90.70 47.40 62.30 25.70 28.50 27.00 44.65
IForest Liu et al. (2008) 56.90 59.80 58.30 43.90 35.30 39.10 48.70
AnomalyTrans Xu et al. (2021) 90.71 47.43 62.29 29.96 48.63 37.08 49.69
DCdetector Yang et al. (2023) 96.57 59.08 73.31 38.41 59.73 46.76 60.04
PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
@@ 4.1 Setups and Evaluation Metrics
We applied eight datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others: github.com/DAMO-DI-ML/KDD2023-DCdetector. Details about these datasets are outlined in Table 4 in Appendix. Also, implementation and pseudo-code Algorithm 1 details in Appendix.
We compared our model with 27 benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; See details in Table 1 and Table 2, or in Appendix Baslines.
Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as accuracy, precision, recall, and F1-score Xu et al. (2021). Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall Huet et al. (2022), along with Volume under the surface (VUS) Paparrizos et al. (2022). Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events. Therefore, we introduce the following two new metrics. Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels. VUS considers the inclusion of anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve. Employing various methodologies allows for a more comprehensive evaluation of our model. To fairly compare with other methods, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
(a) Project constraint
(b) Anomaly ratio
(c) $d_{model}$ size
(d) Encoder layer
(e) Window size
(f) Multi-scale size
@@ 4.2 Comparison Results
Data Source Clarification: Due to substantial device variations, to ensure fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes. The results of AnomalyTrans and DCdetector in Table 1 and Table 2 are our reproductions.
@@
In Table 2, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types. PatchAD consistently achieves SOTA results on these two datasets compared to other baseline methods. Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset.
Given the remarkable performance of AnomalyTrans and DCdetector surpassing other baselines significantly, we conducted a comprehensive comparison of PatchAD with them across all datasets using multiple metrics, as depicted in Table 5 in Appendix. Our proposed method outperforms or at least remains competitive compared to these two SOTA methods. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive.
@@
(c) Time cost in different $d_{model}$
(d) Time cost in different window
@@ 4.3.1 Ablation Studies
To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3. The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes the removal of the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modeling inter-channel information and sharing channel information are beneficial for model learning. Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, ’wo_constraint’ showcases the impact of removing the constraint on the projection head. In conjunction with Figure 3(a), we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
Table 3: Ablation experiments on MSL, SMAP, WADI datasets.
@@
Ours 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
@@ 4.3.2 Parameter Sensitivity
We conducted a sensitivity analysis on PatchAD’s parameters. The default parameter settings remain consistent with the basic configuration. Figure 1(a) demonstrates the impact of the projection head constraint on PatchAD. Overall, it holds more significance for smaller-scale datasets like MSL. However, excessively strong constraints hinder the model’s learning process. Therefore, setting this constraint between 0.1 to 0.4 is more advisable. Figure 1(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2. The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets compared to the MSL dataset. For most datasets, setting this value between 0.8 and 1 yields optimal results. Figure 1(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$. PatchAD demonstrates better results with smaller $d_{model}$ sizes. Therefore, to strike a balance between model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively. Figure 1(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4. Figure 1(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210. The time window is a critical parameter, typically set to an initial value of 105 across most datasets. Lastly, Figure 1(f) displays PatchAD’s performance under different multi-scale patch size combinations. The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
@@ 4.3.3 Visual Analysis
We visualized PatchAD’s response to various anomalies in Figure 3. Synthetic univariate time series were generated by Lai et al. (2021), incorporating anomalies such as global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies. The first row depicts the original data, while the second row displays PatchAD’s output of anomaly scores against the anomaly threshold. In particular, PatchAD has the ability to detect various types of anomalies.

### v5

@@ 3.5 Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both. The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$ (7)
$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
@@
The final PatchAD loss comprises the two aforementioned components and the reconstruction term ($\mathcal{L}_{rec}=\operatorname{MSE}(\hat{\mathcal{X}},\mathcal{X})$), which is accompanied by a constraint coefficient to regulate the strength of the constraint.
$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}+\mathcal{L}_{rec}.$ (10)
@@ 3.6 Anomaly Score
Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best in Bold, and the second in underlined.
@@
OCSVM [19] 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
IForest [37] 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
LOF [8] 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
U-Time [38] 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
ITAD [39] 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
DAGMM [9] 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
VAR [12] 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD [40] 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA [41] 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 [42] 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
BeatGAN [43] 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM-VAE [6] 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BOCPD [44] 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
Deep-SVDD [45] 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
LSTM [46] 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly [7] 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion [47] 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC [48] 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans [27] 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetector [10] 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
TimesNet [14] 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
PatchTST [49] 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
GPT2-Adapter [15] 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
NPSR-pt [50] 10.60 100.0 19.16 24.99 96.37 39.68 27.74 100.0 43.43 12.29 100.0 21.89 31.04
NPSR-seq [50] 17.66 100.0 30.02 12.79 100.0 22.68 37.87 99.95 54.81 86.76 88.58 87.66 48.80
@@ 4.1 Setups and Evaluation Metrics
We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP (3) PSM (4) SMD (5) SWaT (6) WADI (7) NIPS-TS-SWAN (8) NIPSTS-GECCO (9) UCR. Details about these datasets are outlined in Table 4 in Appendix. Also, implementation and pseudo-code Algorithm 1 details in Appendix.
We compared our model with several benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; (8) Large language model; (9) Diffusion model; See details in Table 1 and Table 5, or in Appendix B.2.
Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as AUC, accuracy, precision, recall, F1-score with point adjustment (F1), and original F1-score (F1_ori) [27]. Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall [51], along with Volume under the surface (VUS) [52]. Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events. Therefore, we introduce the following two new metrics. Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels. VUS considers including anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve. Employing various methodologies allows for a more comprehensive evaluation of our model. To compare with other methods fairly, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
@@ 4.2 Comparison Results
Data Source Clarification: Due to substantial device variations, to ensure a fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes. The results of AnomalyTrans and DCdetector in Table 1 and Table 5 are our reproductions.
@@
In Table 5 in Appendix, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types. Compared to other baseline methods, PatchAD consistently achieves SOTA results on these two datasets. Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset. Given the remarkable performance of AnomalyTrans and DCdetector surpassing other baselines significantly, we conducted a comprehensive comparison of PatchAD with them across all datasets using multiple metrics, as depicted in Table 7 in Appendix. Our proposed method outperforms or at least remains competitive compared to these two SOTA methods. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive. Table 2 demonstrates that our approach outperforms previous methods on univariate dataset UCR. This shows that our method can also fetch state-of-the-art performance on univariate datasets.
Table 2: Results on UCR dataset.
Metric Acc P R F1
AnomalyTrans 99.49 60.41 100 73.08
DCdetector 99.51 61.62 100 74.05
PatchAD (Ours) 99.52 72.72 100 84.00
Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (WADI, SWAT, PSM, NIPS-TS-SWAN). PatchAD has demonstrated a remarkable improvement in the F1 score, with a notable increase of 50.5%. Additionally, there has been a 7.8% improvement in the Aff-F1 score and a commendable 10.0% boost in the AUC.
@@
(c) Different $d_{model}$
(d) Different window
@@
4.3 Unveiling the Mechanisms
PatchAD leverages Inter and Intra Mixers to learn detailed and generalized features respectively, as demonstrated in Figure 8. Utilizing multiple patch sizes, as explored in Appendix (Figures 8(a) & 8(b)), enhances feature learning and generalization. The interplay between these mixers, visualized in Figure 3(c), effectively separates and amplifies distinct features, improving anomaly detection. Empirical validation, depicted in Figure 9, reveals that increasing patch size leads to more significant features acquired by the Intra Mixer, accelerating model learning. Finally, Figure 10 shows that anomalies are most easily detected in the Intra Mixer view due to its enhanced distinctiveness. In essence, maximizing the disparity between the Inter and Intra Mixers enhances PatchAD’s detection capabilities while mitigating overfitting. More details are provided in Appendix.
@@ 4.4.1 Ablation Studies
Table 3: Ablation studies on MSL, SMAP, WADI datasets.
@@
PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3. The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes removing the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modelling inter-channel information and sharing channel information benefit model learning. Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, ’wo_constraint’ showcases the impact of removing the constraint on the projection head. In conjunction with Figure 5(a), we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
@@ 4.4.2 Parameter Sensitivity
A sensitivity analysis was conducted to investigate the impact of various PatchAD parameters on model performance (Figure 6). We find that a projection head constraint within the range of 0.1 to 0.4 is generally advisable, particularly for smaller datasets (Figure 6(a)). Optimal anomaly thresholds ($\sigma$) typically fall between 0.8 and 1.0, demonstrating robustness across various datasets (Figure 6(b)). A network dimension ($d_{model}$) of 40 strikes a balance between model complexity and performance (Figure 6(c)). PatchAD exhibits consistent performance across a range of window sizes (Figure 6(e)), and employing 3 or 4 encoder layers is found to be suitable (Figure 6(d)). Finally, combining multiple patch sizes demonstrates a positive impact on overall performance (Figure 6(f)). Detailed analysis for each parameter is provided in Appendix.
@@ 4.4.3 Visual Analysis
We visualized PatchAD’s response to various anomalies in Figure 4. Synthetic univariate time series were generated by [53], incorporating global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies. The first row depicts the original data, while the second row displays PatchAD’s output of anomaly scores against the anomaly threshold. In particular, PatchAD has the ability to detect various types of anomalies. Figure 7 presents the practical performance of PatchAD on two real-world datasets. PatchAD exhibits the capability to detect the majority of anomalies, albeit with two anomalies remaining undetected. Furthermore, a noticeable disparity is observed between the reconstructed features by PatchAD and the original features, indicating its proficiency in learning the underlying patterns within the data, rather than simply replicating the original features.
Figure 7: The case studies on SWAT and WADI datasets.